# What-if Analysis for Business Users: Current Practices and Future Opportunities

### v3

@@ 2.3. Advanced Analytics Tools
Marx et al. (Marx et al., 2012) and Bergeron et al. (Bergeron et al., 1995) studied legacy information systems in organizations or Executive Information Systems (EIS), identifying gaps in decision-making support, especially in drill-down, scenario analysis, and optimization. Van et al. (Van Bruggen et al., 2010) explored systems for marketing management, echoing these limitations. More recent work by Crisan and Correll (Crisan and Correll, 2021) and Bartram et al. (Bartram et al., 2021) also noted these shortcomings, though focused less on business decisions. Zhang et al. (Zhang et al., 2020) and Honeycutt et al. (Honeycutt et al., 2020) emphasized integrating domain knowledge with AI to enhance user trust in intelligent systems. Bhattacharya et al. (Bhattacharya et al., 2024) empowers healthcare workers to utilise AI model explanations to improve model predictions. Oral et al. (Oral et al., 2023) highlight the lack of decision support features like visibility of alternatives, input processing, change model parameters, etc. in existing tools. Our study expands on exploring model-based WIA techniques for making data-driven decisions, highlighting challenges and opportunities for integrating in them domain-specific expertise.
Most human-in-the-loop machine learning research centers on data scientists and analysts, highlighting their openness to automating decisions (Wang et al., 2019) or collaborating with models (Poursabzi-Sangdeh et al., 2021). Crisan and Fiore-Gartland (Crisan and Fiore-Gartland, 2021) discuss automation’s role in routine tasks, rapid prototyping, and democratizing data science, but for data scientists. Our study differs by focusing on how business users with no formal technical backgrounds (hence referred to as non-technical users) perceive advanced WIA, a form of AutoML used for data-driven decisions. Further, we go beyond interviews, collecting empirical data from users actively engaging in hands-on decision-making with these techniques.
Many business intelligence (BI) tools (Ferrari and Russo, 2016; Tableau, 2021b, a; SAS, 2021) and spreadsheet applications offer some advanced analytics, but they primarily focus on descriptive and exploratory tasks, answering the “what” and “why” in data. But, they often lack predictive and prescriptive features, leaving the “now what?” unanswered to plan subsequent actions. Tools like Excel’s SOLVER (Microsoft, 2021a) and GOAL SEEK (Microsoft, 2021b) provide optimization features for desired outputs (Bartram et al., 2021), but manual formula creation remains challenging for business users, which interactive advanced analytics tools aim to address.
@@ 2.4. Organizational Decision-Making
Management and organizational theory categorize decisions into three tiers: strategic (executive level), tactical (middle management), and operational (front-line employees) (Ballou, 1992). This work focuses on tactical decisions. Simon’s three-step model (Simon, 1960) (also observed in (Dimara et al., 2021) and used in (Oral et al., 2023))—intelligence (information gathered and problems identified), design (possible solutions for problems generated and evaluated), and choice (best alternative selected)—describes the decision-making process, as does Berisha-Shaqiri’s eight-step model (Berisha-Shaqiri, 2014)-identification of decision’s purpose, information gathering, principles for judging the alternatives, brainstorm and analyze choices, evaluating alternatives, decision execution, and results evaluation. On similar lines, Jun (Jun, 2022) centers on hypothesis formalization through statistical analyses by data scientists. We observe similar steps in our study participants’ decision-making, but we focus deeper into the data they use, the advanced analytics they apply, and the challenges they face in selecting the best alternatives. Perkins et al. (Perkins and Rao, 1990) examine how marketing managers’ experience influences their decisions and the information they use, while Little (Little, 1979) reports on processes followed by marketing managers from over 40 years ago, but with the aid of descriptive and exploratory tools. In all these works, the role of data and analysis is largely simplified or overlooked. In contrast, our study focuses on how data and advanced analytics are employed by managers across varied departments.
@@ 3. Overview of the Two Studies
This section outlines the methods and procedures for our two-part study.
Participant ID Company Sector Company Size Participant ID Company Sector Company Size P1 Education 10,001+ P12 SaaS 201-1000 P2 Consumer Goods 1001-5000 P13 Human Resources 51-200 P3 Non-Profit 51-200 P14 Healthcare 1001-5000 P4 Telecommunications 201-1000 P15 Consumer Goods 10,001+ P5 E-learning 201-1000 P16 Food 10,001+ P6 SaaS 10,001+ P17 Logistics, Supply Chain 11-50 P7 Health, Wellness, Fitness 5001-10,000 P18 IT 201-1000 P8 Telecommunications 10,001+ P19 Hospitality 10,001+ P9 Construction 5001-10,000 P20 Healthcare 10,001+ P10 Consumer Goods 201-1000 P21 SaaS 10,001+ P11 Electrical/Electronic Manufacturing 5001-10,000 P22 SaaS 201-1000
Table 1. Study participants hailed from four business departments of companies with diverse sizes and sectoral focuses. Colors in the table refer to departments in which the participants worked: marketing, sales, product, and operations.
Figure 2. Overview of our two-part study. We conduct the study with 22 business users (12 female, 10 male) from four departments: marketing (7), sales (5), product (5), and operations (5). The interview study aims to understand various what-if analysis (WIA) techniques participants employ, the tools they use, and the challenges they face to use them for their data-driven decision-making of both their own business use cases as well as for a specific business use case of Marketing Mix Modeling (MMM). The task-based study aims to understand the same participant’s perception of practically using representative WIA techniques identified in the first study and implemented as a probe within a visual data analytics prototype. Colors refer to the departments in which the participant worked (marketing, sales, product, operations).
@@ 3.1. Participants
We recruited 22 business users through an online survey (Anonymous, 2024b) through the User Interviews (Interviewss, 2021) website and company’s mailing lists. Our survey ensured that participants were selected based on three criteria: (1) use organizational, customer, or product data to guide business decisions, (2) make decisions at least weekly, and (3) have no formal training in data science or programming. Final participant pool comprised business users from four departments: marketing (7), sales (5), product (5), and operations (5) managers, from varied company sizes, sectors, and expertise levels, as shown in Table 1.
@@ 3.2. Protocol
Each participant completed both parts of the study over 120–140 minutes in two sessions, held on separate days within one to two weeks, as illustrated in Figure 2. All interviews were conducted and recorded online for analysis. Participants received a total of \$175 in Amazon gift card for completing both sessions. Each study is detailed in later sections.
@@ 3.3. Analysis
We analyzed the recorded and transcribed data from both parts of the study using iterative coding to derive our findings. In the first interview study, a researcher initially categorized participants’ business decisions, data usage, and general WIA techniques and tools. The coding then focused on specific WIA techniques, such as scenario exploration and ranking correlations between drivers and key performance indicators (KPIs) among many others, to understand their usage and challenges. Finally, both these analyses were then synthesized to form the overall findings of the interview study.
In the follow-up task-based study, participants’ experiences of practically using four WIA techniques identified from the previous study and implemented within a visual analytics prototype as a probe are captured. Their feedback on potential benefits and improvement opportunities of WIA techniques were similarly categorized. As new data emerged, these categories were refined, with any ambiguities addressed through team discussions. We support our findings with representative participant quotes throughout the paper.
@@ 4. Study I: Interview Study
We present the method and findings of the first interview study.
@@ 4.1. Method
The interview study was conducted through semi-structured interviews with business users to understand various WIA techniques they employed, motivations behind their use, tools they relied on, and challenges they encountered. First, we describe our method as outlined in Figure 2.
@@ 4.1.1. Protocol
General Session. The general session involved 21 open-ended questions to establish the broader context about their decision-making and lasted 45 minutes. We began by exploring the types of decisions participants make, the data they use, and the teams they rely on. Then, we transitioned to discussing their general advanced analytics techniques they use, such as WIA or predictive and prescriptive model-based analytics. Overall questions were grouped into four categories: (1) role and team, (2) business goals, (3) data, tools, and WIA techniques, and (4) ideal tools, as detailed in Table 2.
# Interview Study Questionnaire Role and Team 1 Briefly describe your role in your company? How many people are part of your team and what are their roles? 2 Are there data analysts in your team? Business Goals 3 How often does your daily work require you to answer questions or make decisions to achieve your business objectives? 4 Describe recent business questions that you/your team often try to answer. 5 Approximately how many hours or days do you spend answering each of these questions? 6 Do you answer these questions in one analysis sitting or repeatedly come back to find answers that would eventually help make a decision for your team or company? Data, Tools, and WIA Techniques 7 How often do you/your team use the data to make decisions? 8 Describe the data you or your team utilize in your work. 9 What are the tools or systems (could be commercial, your company products, or self-developed) you generally use to answer such questions? 10 Where does this data come from? 11 Talk us through the analytic process of coming up with strategies or making decisions elaborating on analytic techniques you use in response to one or two specific business questions you shared? 12 Do you consider your current process of using analytics you use efficient? Explain. 13 Do the set of tools you use for performing the analytics you use seem efficient? 14 What are the challenges you face to achieve your objectives? 15 What are the challenges of the set of tools you use? 16 Do you use any advanced analytics like making predictions using the data for your decision-making? Why/Why not? 17 Do you base your predictions on your intuition? Explain. 18 Do you base your predictions on your experience? Explain. 19 Have you ever heard of WIA, predictive analytics, and prescriptive analytics? Can you describe some examples? 20 Do you use any of these analytics that use models or similar techniques (e.g., machine learning, deep learning, statistical modeling, techniques, etc.)? If yes, which ones, why, which tools, expertise level, number of years of experience, examples of how. If no, why not and which techniques would you like to use? If sometimes, why? Ideal Tool 21 If you were to invent the perfect tool to work with data and make decisions with, what would that tool look like? What would it do?
@@
Table 2. Questions that we asked participants during the general session of the interview study. These questions aimed to understand business users’ roles and teams, business goals, data, tools, WIA practices, and desired tools for making data-driven decisions.
@@
Task-Based Session. After the general session, participants spent 15-25 minutes on a task-based session discussing WIA techniques, tools, and challenges in a specific business use case of Marketing Mix Modeling (MMM). Discussion on this specific use case in addition to the general discussion allowed for consistent comparisons across participants.
@@
Participants were shown this dataset, sourced from Kaggle (Kaggle, 2022) in CSV format during task briefing. And prompted to discuss advanced analysis techniques and tools they would employ to analyze the data and make decisions to maximize sales. This use case was chosen since it was relatable for many, especially marketing managers, and broad enough for non-marketing participants to apply their own experiences to. Assumptions and validity of both this use case and dataset were verified by our company’s marketing experts, who served as our collaborators. Participants could take notes, if desired.
@@ 5. Findings of the Interview Study
We present the interview study findings in three parts: participants’ decision-making context (goals, decisions, and data used), the WIA techniques and tools employed along with specific challenges encountered, and the broader challenges associated with advanced analytics.
5.1. Business Users’ Goals and Decisions
@@
Business User KPI Goals Decision Questions Data Marketing click rate, inquiry rate, conversion rate, retention rate which [existing running] campaigns are doing well and why? (P1, P13), how to optimize media mix? (P19), how to move opportunities higher in sales [pipeline]? (P1, P8) search engine optimization (SEO) data (e.g, traffic, impressions, page rank), click rate, customer usage and engagement data, budget data, sales pipeline data, e-mail activity Sales conversion rate, retention rate, churn rate, return of investment what actions must be taken to increase customer retention rate? (P22), how can we leverage previous or present customer to increase sales now? (P9, P14, P18), how do ecological conditions affect churn behavior in comparison to internal factors? (P5) sales pipeline data, revenue and sales data, quote data, product usage data, customer success scores and metrics, campaign data, and external data like shipping costs and delays, geographical data, competitor data Product CSAT score, retention rate, conversion rate, click rate, content management which features of the product (pricing, audience, etc.) and product changes/enhancements are needed in competing markets or for a given geography? (P4, P12), which customer verticals [or segment of a marketplace] are effective? how to increase their market share? how to tap the others? (P2, P4), should we invest in customer needs given the risk? (P2, P11) sales data, market share data, customer success data, CRM data, customer demographics data, customer interaction and engagement data, customer activities, feature usage Operations CSAT score, throughput, productivity, inventory stock how to handle staffing [hiring/management] to reduce costs, maintain productivity, and handle ecological condition changes (e.g., holiday season, inflation, etc.)? (P6, P15), how to cut down wait-times for both products and customers? (P16), how to optimize inventory efficiently? (P16, P17) inventory data (number of products, types of products, count of products, etc.), budget and pricing data, CSAT data, historical data of buy, sell, profits, and sales made from various products, manufacturing or development time, processing time, billing data, time aspects, failure or success data
Table 3. Goals, decision questions, and data sources considered by study participants when making decisions at work.
@@ 5.1.2. Decisions
To achieve their goals, business users made decisions about actions related to drivers, which were data variables about customers and products. Examples included search engine optimization (SEO) data (e.g., traffic, impressions), product usage data (e.g., feature usage, time spent), and customer activities (e.g., sign-ups, clicks). For instance, to boost conversion rates, decisions included “providing 3 months free trial on the product on signing up for a demo” (which involved the “demo sign up” driver), or “sending out targeted marketing emails showing how the product works for prospective customers’ use cases” (which involved the “marketing emails sent” driver). Similarly, to increase retention rate, plausible decision was “changing placements of product features and highlighting certain features using pop-ups that match customers in a certain geography” (which involved the “feature usage frequency” driver), or “providing tutorial use cases of product mapped to customers’ business niche” (which involved the “customer vertical” driver). More examples are tabulated in Table 3. For a given goal, participants needed to decide both which drivers to act on and how to modify them.
5.2. WIA Techniques Employed by Business Users
Participants used various advanced analytics techniques to understand how independent driver variables relate to dependent KPI business goals and to form hypotheses about these driver variables as decision actions.

### v4

@@ 2.3. Advanced Analytics Tools
Marx et al. (Marx et al., 2012) and Bergeron et al. (Bergeron et al., 1995) studied legacy information systems in organizations or Executive Information Systems (EIS), identifying gaps in decision-making support, especially in drill-down, scenario analysis, and optimization. Van et al. (Van Bruggen et al., 2010) explored systems for marketing management, echoing these limitations. More recent work by Crisan and Correll (Crisan and Correll, 2021) and Bartram et al. (Bartram et al., 2021) also noted these shortcomings, though they focused less on business decisions. Zhang et al. (Zhang et al., 2020) and Honeycutt et al. (Honeycutt et al., 2020) emphasized integrating domain knowledge with AI to enhance user trust in intelligent systems. Bhattacharya et al. (Bhattacharya et al., 2024) empower healthcare workers to utilize AI model explanations to improve model predictions. Oral et al. (Oral et al., 2023) highlight the lack of decision support features like the visibility of alternatives, input processing, change model parameters, etc. in existing tools. Our study expands on exploring model-based WIA techniques for making data-driven decisions, highlighting challenges and opportunities for integrating domain-specific expertise into them.
Most human-in-the-loop machine learning research centers on data scientists and analysts, highlighting their openness to automating decisions (Wang et al., 2019) or collaborating with models (Poursabzi-Sangdeh et al., 2021). Crisan and Fiore-Gartland (Crisan and Fiore-Gartland, 2021) discuss automation’s role in routine tasks, rapid prototyping, and democratizing data science, but for data scientists. Dimara et al. (Dimara et al., 2021) share similar decision-making practices of decision makers. Our study differs by focusing on how specific group of users; business professionals who perform analysis using non-technical means perceive advanced WIA, a form of AutoML used for data-driven decisions. Further, we go beyond interviews, collecting empirical data from users actively engaging in hands-on decision-making with these techniques.
Many business intelligence (BI) tools (Ferrari and Russo, 2016; Tableau, 2021b, a; SAS, 2021) and spreadsheet applications offer some advanced analytics, but they primarily focus on descriptive and exploratory tasks, answering the “what” and “why” in data. However, they often lack predictive and prescriptive features, leaving the question of “now what?” unanswered, making it difficult to plan subsequent actions. Tools like Excel’s SOLVER (Microsoft, 2021a) and GOAL SEEK (Microsoft, 2021b) provide optimization features for desired outputs (Bartram et al., 2021), but manual formula creation remains challenging for business professionals, which interactive advanced analytics tools aim to address.
@@ 2.4. Organizational Decision-Making
Management and organizational theory categorize decisions into three tiers: strategic (executive level), tactical (middle management), and operational (front-line employees) (Ballou, 1992). This work focuses on tactical decisions, which are short-to-medium term actionable steps (e.g., introduce discounts, run campaigns, reduce shipping costs, etc.) aimed at achieving specific business goals (e.g., compete with competitors’ price changes, decrease churn rate, increase throughput, etc.). Additional examples are provided in Section 5.1. These decisions are dynamic, continuously evolving, and need improvement as they are central to the responsibilities of business professionals.
The three-step model by Simon (Simon, 1960) (also observed in (Dimara et al., 2021) and used in (Oral et al., 2023))—intelligence (information gathered and problems identified), design (possible solutions for problems generated and evaluated), and choice (best alternative selected)—describes the decision-making process. Berisha-Shaqiri provides a more detailed eight-step model (Berisha-Shaqiri, 2014). On similar lines, Jun (Jun, 2022) centers on hypothesis formalization through statistical analyses by data scientists. We observe similar steps in our participants’ decision-making, but we focus on the data they use, the advanced analytics they apply, and the challenges they face in selecting the best alternatives. Perkins et al. (Perkins and Rao, 1990) examine how marketing managers’ experience influences their decisions and the information they use, while Little (Little, 1979) reports on processes followed by marketing managers from 40 years ago, but with the aid of descriptive and exploratory tools. In all these works, the role of data and analysis is largely simplified or overlooked. In contrast, our study focuses on how managers employ data and advanced analytics across varied departments.
@@ 3. Overview of the Two Studies
Here, we outline the methods and procedures for our studies.
Figure 1. Overview of our two-part study. We conduct the study with 22 business professionals (12 female, 10 male) from four departments: marketing (7), sales (5), product (5), and operations (5). The interview study aims to understand various what-if analysis (WIA) techniques participants employ, the tools they use, and the challenges they face to use them for their data-driven decision-making of both their own business use cases as well as for a specific business use case of Marketing Mix Modeling (MMM). The task-based study aims to understand the same participant’s perception of practically using representative WIA techniques identified in the first study and implemented as a probe within a visual data analytics prototype. Colors refer to the departments in which the participant worked (marketing, sales, product, operations).
[…]
[Overview of our two-part study.]Overview of our two-part study. We conduct the study with 22 business professionals (12 female, 10 male) from four departments: marketing (7), sales (5), product (5), and operations (5). The interview study aims to understand various what-if analysis (WIA) techniques participants employ, the tools they use, and the challenges they face to use them for their data-driven decision-making of both their own business use cases as well as for a specific business use case of Marketing Mix Modeling (MMM). The task-based study aims to understand the same participant’s perception of practically using representative WIA techniques identified in the first study and implemented as a probe within a visual data analytics prototype.
@@ 3.1. Participants
We recruited 22 business professionals via an online survey (Anonymous, 2024b) through the User Interviews (Interviewss, 2021) website and company’s mailing lists. Our survey ensured that participants were selected based on three criteria: (1) use organizational, customer, or product data to guide business decisions, (2) make decisions at least weekly, and (3) have no formal training in data science or programming. The final participant pool comprised of business professionals from four departments: marketing (7), sales (5), product (5), and operations (5) managers from varied company sizes, sectors, and expertise levels, as shown in Table 1 in supplementary materials.
@@ 3.2. Protocol
Each participant completed both parts of the study in 120–140 minutes in two sessions, held on separate days within one to two weeks, as illustrated in Figure 1. All interviews were conducted, recorded, and transcribed online for analysis. Participants received a total of \$175 in Amazon gift card for completing both sessions. Each study is detailed in later sections.
@@ 3.3. Analysis
We analyzed the recorded and transcribed data from both parts of the study using iterative coding to derive our findings. The first author, hereafter referred to as the coder, primarily conducted the coding. The coder double-checked with the other co-authors whenever questions arose or clarifications were needed to ensure rigor and consistency. For the first interview study, the coder began by conducting an initial broad categorization on participants’ business decisions, data utilization strategies, and general WIA techniques and tools. This included identifying recurring patterns in how participants approached making data-driven decisions and integrated WIA into their workflows. In the second iteration, the coder focused on understanding the usage of specific WIA techniques (e.g., scenario exploration and ranking correlations between drivers and key performance indicators (KPIs)) and associated challenges (e.g., current tools being simplistic, needing manual efforts). In the third iteration, results from the first two rounds were then synthesized to form the overall findings of the interview study, providing a comprehensive understanding of participants’ WIA workflows and challenges when making data-driven decisions.
For the follow-up task-based study, the coder analyzed participants’ experiences of practically using four WIA techniques identified from the previous interview study and implemented within a visual analytics prototype probe. Over subsequent iterations, the analysis was revised to categorize the use of the functionalities, immediate reactions, challenges, and feedback. The coder then categorized these findings further, grouping feedback into potential benefits (Section 7.2) and opportunities (Section 7.3).
In both studies, the coder consulted with the other authors over subsequent iterations to ensure validity. This collaborative process ensured that the categories were coherent and accurately reflected the data. Ambiguities, such as conflicting interpretations of a participant’s comments, were resolved through discussions with all authors. We support our findings with representative participant quotes throughout the paper.
@@ 4. Study I: Interview Study
Here, we present the method of the first interview study.
@@ 4.1. Method
The interview study was conducted through semi-structured interviews with business professionals to understand various WIA techniques they employed, motivations behind their use, tools they relied on, and challenges they encountered. It is outlined in Figure 1 and described underneath.
@@ 4.1.1. Protocol
General Session. The general session involved 21 open-ended questions to establish the broader context about their decision-making. These open-ended questions enabled participants to share their existing workflows in-depth and flexibly using their own use cases, offering insights across four categories: (1) role and team, (2) business goals, (3) data, tools, and WIA techniques, and (4) ideal tools, as detailed in Table 2. This session lasted 45 minutes.
# Interview Study Questionnaire
Role and Team
1 Briefly describe your role in your company. How many people are on your team, and what are their roles?
2 Are there data analysts in your team?
Business Goals
3 How often does your daily work require you to answer questions or make decisions to achieve your business objectives?
4 Describe recent business questions that you/your team often try to answer.
5 Approximately how many hours or days do you spend answering each of these questions?
6 Do you answer these questions in one analysis sitting or repeatedly come back to find answers that would eventually help make a decision for your team or company?
Data, Tools, and WIA Techniques
7 How often do you/your team use the data to make decisions?
8 Describe the data you or your team utilize in your work.
9 What are the tools or systems (could be commercial, your company products, or self-developed) you generally use to answer such questions?
10 Where does this data come from?
11 Talk us through the process of coming up with strategies or making decisions, elaborating on analytic techniques you use in response to one or two specific business questions you shared.
12 Do you consider your current process of using analytics you use efficient? Explain.
13 Do the set of tools you use for performing your analytics seem efficient?
14 What are the challenges you face to achieve your objectives?
15 What are the challenges of the set of tools you use?
16 Do you use any advanced analytics like making predictions using the data for your decision-making? Why/Why not?
17 Do you base your predictions on your intuition? Explain.
18 Do you base your predictions on your experience? Explain.
19 Have you ever heard of WIA, predictive analytics, and prescriptive analytics? Can you describe some examples?
20 Do you use any of these analytics that use models or similar techniques (e.g., machine learning, deep learning, statistical modeling, techniques, etc.)?
If yes, which ones, why, which tools, expertise level, number of years of experience, examples of how. If no, why not and which techniques would you like to use? If sometimes, why?
Ideal Tool
21 If you were to invent the perfect tool to work with data and make decisions with, what would that tool look like? What would it do?
@@ 15 What are the challenges of the set of tools you use?
Table 2. Questions that we asked participants during the general session of the interview study. These questions aimed to understand the context around business professionals’ roles and teams, business goals, data, tools, WIA practices, and desired tools for making data-driven decisions.
@@
Task-Based Session. Participants then spent 15-25 minutes on a task-based session discussing WIA techniques, tools, and challenges for a specific business use case of Marketing Mix Modeling (MMM). Discussion on this specific use case allowed for consistent comparisons across participants.
@@
Participants were shown this dataset, sourced from Kaggle (Kaggle, 2022) in a CSV format during task briefing. And prompted to discuss advanced analysis techniques and tools they would employ to analyze the data and make decisions to maximize sales. This use case was chosen since it was relatable for many, especially marketing managers, and broad enough for non-marketing participants to apply their own experiences to as well. Assumptions and validity of both this use case and dataset were verified by marketing experts at a company, who served as our collaborators. Participants could take notes.
@@ 5. Findings of the Interview Study
We present the interview study findings in three parts: participants’ decision-making context (goals, decisions, and data used), the WIA techniques and tools employed along with specific challenges encountered, and the broader challenges with advanced analytics.
5.1. Business Professionals’ Goals and Decisions
@@ 5.1.2. Decisions
To achieve their goals, business professionals made decisions about actions related to drivers, which were data variables about customers and products. Examples included search engine optimization (SEO) data (e.g., traffic, impressions), product usage data (e.g., feature usage, time spent), and customer activities (e.g., sign-ups, clicks). For instance, to boost conversion rates, decisions included “providing 3 months free trial on the product on signing up for a demo” (which involved the “demo sign up” driver), or “sending out targeted marketing emails showing how the product works for prospective customers’ use cases” (which involved the “marketing emails sent” driver). Similarly, to increase retention rate, a plausible decision was “changing placements of product features and highlighting certain features using pop-ups that match customers in a certain geography” (which involved the “feature usage frequency” driver), or “providing tutorial use cases of product mapped to customers’ business niche” (which involved the “customer vertical” driver). More examples are included in Table 2 in supplementary materials. For a given goal, participants needed to decide both which drivers to act on and how to modify them.
5.2. WIA Techniques Employed by Business Professionals
Participants used various advanced analytics techniques to understand how independent driver variables relate to dependent KPI business goals and to form hypotheses about these driver variables.